# Null projections and noncommutative function theory in operator algebras

### v1

@@ 3. Measurably indistinguishable projections
Since $\{t\}$ has measure zero, we see that $\lambda(p_{t})=0$ so that $p_{t}\leq 1-e$ for each $t\in{\mathbb{T}}$, and thus $\bigvee_{t\in{\mathbb{T}}}p_{t}\leq 1-e$. Fix a Borel measurable subset $E\subset{\mathbb{T}}$ with non-zero measure. We clearly have $\bigvee_{t\in E}p_{t}\leq\chi_{E}$. On the other hand, $\chi_{E}(\lambda)>0$ so $\chi_{E}$ is not dominated by $1-e$, and in particular, $\chi_{E}\neq\bigvee_{t\in E}p_{t}$.
@@
Let $B$ be a $\mathrm{C}^{*}$-algebra. By considering the usual universal representation of $B$, we find a Hilbert space $H_{u}$ such that $B^{**}\subset B(H_{u})$. We let $(B_{\text{sa}})^{m}\subset B^{**}$ denote the set of elements in $B^{**}$ that can be obtained as the increasing limit in the SOT of $B(H_{u})$ of a self-adjoint net from $B$. A self-adjoint element $x\in B^{**}$ is said to be universally measurable if, for every $\varepsilon>0$ and state $\varphi$ on $B$, there are elements $y,z\in(B_{\text{sa}})^{m}$ such that $-y\leq x\leq z$ and $\varphi(y+z)<\varepsilon$.
Projections in $B^{**}$ that are either open or closed are universally measurable [pedersen1979book, Proposition 3.11.9]. More generally, the class of universally measurable projections also contains the Borel projections in the sense of [combes1970]; see [brown1970] for more detail.
@@
Let $z_{\operatorname{at}}\in B^{**}$ denote the supremum of all minimal projections in $B^{**}$. It is well known that $z_{\operatorname{at}}$ is central: indeed, it coincides with the support projection for the unique weak-*** continuous extension to $B^{**}$ of the atomic representation of $B$ [pedersen1979book, Paragraph 4.3.7]. In particular, by virtue of [pedersen1979book, Lemma 4.3.8], we see that there is a collection $\{H_{\lambda}:\lambda\in\Lambda\}$ of Hilbert spaces such that $z_{\operatorname{at}}B^{**}\cong\prod_{\lambda\in\Lambda}B(H_{\lambda})$. Consequently, if $p\in B^{**}$ is a projection, we have
@@
Let $p,e\in B^{**}$ be universally measurable projections with $e\geq pz_{\operatorname{at}}$. Since $z_{\operatorname{at}}$ is central, we find $(e-p)z_{\operatorname{at}}\geq 0$. In turn, because $e-p$ is universally measurable [pedersen1979book, Proposition 4.3.13], we infer that $e\geq p$ by virtue of [pedersen1979book, Theorem 4.3.15]. ∎
@@ 4. The non-commutative Varopoulos phenomenon
Let $B$ be a $\mathrm{C}^{*}$-algebra, let $q\in B^{**}$ be a closed projection and let $n\geq 1$. Let $R\in{\mathbb{M}}_{1,n}(qB^{**})$ and $C\in{\mathbb{M}}_{n,1}(B^{**}q)$. Fix $\Delta\in{\mathbb{M}}_{n}(B)$, a functional $\varphi\in B^{*}$ and a number $\varepsilon>0$. Then, there are $S\in{\mathbb{M}}_{1,n}(qB)$, $T\in{\mathbb{M}}_{n,1}(Bq)$ such that $\|S\|\leq\|R\|$,$\|T\|\leq\|C\|$ and $|\varphi(R\Delta C-S\Delta T)|<\varepsilon.$
@@
By [hay2007, Proposition 3.1], it follows that $\Phi$ is a completely isometric isomorphism. Hence, $\Phi^{**}:(B/J)^{**}\to(qB)^{**}$ is a completely isometric weak-*** homeomorphism. In turn, by [BLM2004, Paragraph 1.4.4], there is a completely isometric weak-*** homeomorphism $\Psi:(B/J)^{**}\to B^{**}/J^{\perp\perp}$ such that
@@
Next, since $q$ is closed, we obtain that $J^{\perp\perp}=(1-q)B^{**}$. Hence, the map $\Theta:B^{**}/J^{\perp\perp}\to qB^{**}$ defined as
@@
is also a completely isometric weak-*** homeomorphism. Therefore, we obtain a completely isometric weak-*** homeomorphism $\Omega=\Theta\circ\Psi\circ(\Phi^{**})^{-1}:(qB)^{**}\to qB^{**}$ acting as the inclusion on $qB$.
@@
acting as the inclusion on ${\mathbb{M}}_{1,n}(qB)$. Then, the map $\Xi=\Omega^{(n)}\circ\Gamma:{\mathbb{M}}_{1,n}(qB)^{**}\to{\mathbb{M}}_{1,n}(qB^{**})$ is an isometric weak-*** homeomorphism acting as the inclusion on ${\mathbb{M}}_{1,n}(qB)$. Consider the weak-*** continuous linear functional $\psi$ on ${\mathbb{M}}_{1,n}(qB)^{**}$ defined as
@@
By Goldstine’s lemma, we can find $S_{0}\in{\mathbb{M}}_{1,n}(qB)$ with $\|S_{0}\|\leq\|\Xi^{-1}(R)\|=\|R\|$ such that
@@
To find $T$, we argue in a similar fashion. First, proceed as above to find an isometric weak-*** homeomorphism $\Xi^{\prime}:{\mathbb{M}}_{n,1}(Bq)^{**}\to{\mathbb{M}}_{n,1}(B^{**}q)$ acting as the inclusion on ${\mathbb{M}}_{n,1}(Bq)$. Consider the weak-*** continuous linear functional $\psi^{\prime}$ on ${\mathbb{M}}_{n,1}(Bq)^{**}$ defined as
@@
Let $B$ be a $\mathrm{C}^{*}$-algebra and let $A\subset B$ be a subalgebra. A projection $e\in B^{**}$ is said to be $A$-invariant if $ae=eae$ for each $a\in A$. Further, $e$ is said to be $A$-coinvariant if $1-e$ is $A$-invariant. We can now state and prove one of the main results of the paper.
@@
Let $10<\varepsilon<$ 1 and let $\varphi\in A^{\perp}$. We may assume without loss of generality that $\|\varphi\|=1$, so that $|\varphi|$ is a state. Throughout, we put $\tau=|\varphi|$. There is a partial isometry $y\in B^{**}$ such that $\tau=\varphi(\cdot y)$ and $\varphi=\tau(\cdot y^{*})$. By Goldstine’s lemma, there is $t\in B$ such that $\|t\|\leq 1$ and $|\varphi(qy-qt)|<\varepsilon.$ Because $q$ is a complete $(A,\kappa)$-interpolation projection, there is $x\in A$ such that $qx=qt$ and $\|x\|\leq\kappa$. In particular, we have
(5) $|\tau(q)|=|\varphi(qy)|<\varepsilon+|\varphi(qx)|.$
@@ 4.1. Atomic parts and double annihilators
Our class of examples will involve tensor products. We recall some relevant facts here, since several preliminaries are needed. Let $B$ and $D$ be $\mathrm{C}^{*}$-algebras such that $B$ is nuclear. It follows from [AB1980, Theorem 3.1] that there is an injective ***-homomorphism $\sigma:B^{**}\otimes_{\min}D^{**}\to(B\otimes_{\min}D)^{**}$ with the following properties.
@@
    Let $\sigma_{B}:B^{**}\to(B\otimes_{\min}D)^{**}$ and $\sigma_{D}:D^{**}\to(B\otimes_{\min}D)^{**}$ be the restriction ***-homomorphisms arising from $\sigma$ (see for instance [BO2008, Theorem 3.2.6]). Then, $\sigma_{B}$ and $\sigma_{D}$ are both weak-*** continuous.
@@
We first assume that $r$ is open. Let $(d_{j})$ be a net of positive contractions in $D$ increasing to $r$ in the weak-*** topology of $D^{**}$. Then $(\sigma_{D}(d_{j}))$ increases to $\sigma_{D}(r)$ in the weak-*** topology of $(B\otimes_{\min}D)^{**}$. On the other hand, by Goldstine’s theorem there is a net $(a_{i})$ of contractions in $A$ converging to $q$ in the weak-*** topology of $B^{**}$. We see that the net $(\sigma_{B}(a_{i})\sigma_{D}(d_{j}))_{(i,j)}$ converges to $\sigma(q\otimes r)$ in the weak-*** topology of $(B\otimes_{\min}D)^{**}$, being the product of a strongly converging and a weak-*** converging net. Each $\sigma_{B}(a_{i})\sigma_{D}(d_{j})=\sigma(a_{i}\otimes d_{j})$ lies in the image of $A\otimes D$ inside of $(B\otimes_{\min}D)^{**}$, and therefore $q\otimes r\in(A\otimes_{\min}D)^{\perp\perp}$.
@@
Let $q\in(\mathrm{C}(X)\otimes_{\min}D)^{**}$ be a minimal projection. By Lemma 4.5, there are minimal projections $e\in\mathrm{C}(X)^{**}$ and $f\in D^{**}$ such that $q=e\otimes f$. On the other hand, by [gamelin1969, Theorem II.11.3], we see that every $x\in X$ is an $A$-peak point. In other words, every minimal projection in $\mathrm{C}(X)^{**}$ lies in $A^{\perp\perp}$. In particular, this means that $e\in A^{\perp\perp}$, so by Lemma 4.4, we conclude that $p\in(A\otimes_{\min}D)^{\perp\perp}$. This proves the first statement. The second statement immediately follows from the first along with (2). ∎
@@ 4.2. Recovering Varopoulos’ null projections
    there is a net $(b_{i})$ of positive contractions in $B$ commuting with $A$ and decreasing to $q$ in the weak-*** topology of $B^{**}$,
@@ 5.1. Adapted projections
Let $B$ be a $\mathrm{C}^{*}$-algebra and let $A\subset B$ be a subalgebra. We say that a projection $e\in B^{**}$ is $A$-adapted if there is a net $(e_{i})$ of closed $A$-coinvariant projections in $B^{**}$ increasing to $e$ in the weak-*** topology of $B^{**}$. The linchpin technical result of this section is the following refinement of Theorem 4.3.
@@
Let $B$ be a unital $\mathrm{C}^{*}$-algebra and let $A\subset B$ be a unital norm-closed subalgebra. Assume that there is a central closed projection $z\in B^{**}$ such that $B^{**}z$ is commutative and $1-z$ is $A$-adapted. Let $q\in B^{**}$ be a projection with the following properties:
@@
Further, because $1-z$ is $A$-adapted, there is an $A$-coinvariant closed projection $r\leq 1-z$ such that
@@
Because $z$ is closed, so is $z(1-u)$ [akemann1969, Proposition II.5]. Moreover, we have $r\leq 1-z$ so that $r^{\prime}=r+z(1-u)$ is closed by [akemann1969, Theorem II.7]. The projection
@@
Observe next that since $B^{**}z$ is commutative, $q=qz$ is $A$-coinvariant and $uz$ is $A$-invariant. On the other hand, $1-z$ is central and $r$ is $A$-coinvariant, so that $(1-z)-r$ is $A$-invariant. It follows that $u^{\prime}$ is $A$-invariant as well. By virtue of Theorem 4.3, we see that
@@ 5.2. Hilbert function spaces and their multipliers
A multiplier is a function $\varphi:{\mathbb{B}}_{d}\to{\mathbb{C}}$ satisfying $\varphi f\in{\mathcal{H}}$ for every $f\in{\mathcal{H}}$. The corresponding multiplication operator $M_{\varphi}:{\mathcal{H}}\to{\mathcal{H}}$ is then bounded. The space of all multiplication operators induced by a multiplier is denoted by ${\mathcal{M}}({\mathcal{H}})$, and it is weak-*** closed subalgebra of $B({\mathcal{H}})$. If $\varphi$ is a multiplier, then
@@
For each finite subset $F\subset{\mathbb{B}}_{d}$, let $e_{F}\in B({\mathcal{H}})$ denote the projection onto the subspace spanned by $\{k_{\lambda}:\lambda\in F\}$. Since the set $\{k_{\lambda}:\lambda\in{\mathbb{B}}_{d}\}$ spans a norm-dense subset of ${\mathcal{H}}$, the net $(e_{F})$ increases to the identity on ${\mathcal{H}}$ in the weak-*** topology of $B({\mathcal{H}})$. Further, each $e_{F}$ is compact on ${\mathcal{H}}$, and it is ${\mathcal{A}}({\mathcal{H}})$-coinvariant by virtue of (16). Next, by [CT2023Henkin, Lemma 5.5], there is a weak-*** homeomorphic ***-isomorphism $\Omega:{\mathcal{T}}({\mathcal{H}})^{**}(1-z_{K})\to B({\mathcal{H}})$ such that $\Omega(t(1-z_{K}))=t$ for every $t\in{\mathcal{T}}({\mathcal{H}})$. Since $e_{F}$ is compact, it lies in ${\mathcal{T}}({\mathcal{H}})$ and we see that $\Omega^{-1}(e_{F})=e_{F}(1-z_{K})=e_{F}$. It follows that $(\Omega^{-1}(e_{F}))$ is an increasing net of ${\mathcal{A}}({\mathcal{H}})$-coinvariant closed projections in ${\mathcal{T}}({\mathcal{H}})^{**}$ increasing to $1-z_{K}$ in the weak-*** topology of ${\mathcal{T}}({\mathcal{H}})^{**}$. ∎
Let $\gamma:{\mathcal{T}}({\mathcal{H}})\to{\mathcal{T}}({\mathcal{H}})/K$ denote the quotient map. There is a natural ***-isomorphism $\tau:{\mathcal{T}}({\mathcal{H}})/K\to\mathrm{C}({\mathbb{S}}_{d})$ such that
@@
See e.g. [GHX04]. In particular, ${\mathcal{T}}({\mathcal{H}})$ is nuclear [BO2008, Exercise 3.8.1]. The restriction of $(\tau\circ\gamma)^{**}$ gives a weak-*** homeomorphic ***-isomorphism $\Theta:{\mathcal{T}}({\mathcal{H}})^{**}z_{K}\to\mathrm{C}({\mathbb{S}}_{d})^{**}$. We now give a description of the closed projections in ${\mathcal{T}}({\mathcal{H}})^{**}$ dominated by $z_{K}$.
@@
(i) Let $(t_{i})$ be a net of positive contractions in ${\mathcal{T}}({\mathcal{H}})$ decreasing to $q$ in the weak-*** topology of ${\mathcal{T}}({\mathcal{H}})^{**}$. Since $\Theta$ is a weak-*** continuous ***-homomorphism with $\Theta(z_{K})=1$, it follows that $(\Theta(t_{i}z_{K}))$ is a net of positive contractions in $\mathrm{C}({\mathbb{S}}_{d})$ decreasing to $\Theta(q)$ in the weak-*** topology of $\mathrm{C}({\mathbb{S}}_{d})^{**}$. In other words, $\Theta(q)$ is closed in $\mathrm{C}({\mathbb{S}}_{d})^{**}$, so there is a closed subset $E\subset{\mathbb{S}}_{d}$ such that $\Theta(q)=\chi_{E}$.
@@
$\{\varphi:M_{\varphi}\in{\mathcal{A}}({\mathcal{H}})\}\,\chi_{E}=\mathrm{C}({\mathbb{S}}_{d})\,\chi_{E}$
@@
    ${\mathcal{A}}({\mathcal{H}})q={\mathcal{T}}({\mathcal{H}})q$.
@@
By Lemma 5.3, there is a closed subset $E\subset{\mathbb{S}}_{d}$ such that $\Theta(q)=\chi_{E}$. Since $\Theta$ is a ***-isomorphism such that

### v2

@@ 3. Measurably indistinguishable projections
Since $\{t\}$ has measure zero, we see that $\lambda(p_{t})=0$ so that $p_{t}\leq 1-e$ for each $t\in{\mathbb{T}}$, and thus $\bigvee_{t\in{\mathbb{T}}}p_{t}\leq 1-e$. Fix a Borel measurable subset $E\subset{\mathbb{T}}$ with non-zero measure. We clearly have $\bigvee_{t\in E}p_{t}\leq\chi_{E}$. On the other hand, $\chi_{E}(\lambda)>0$ so $\chi_{E}$ is not dominated by 1e1-e, and in particular, $\chi_{E}\neq\bigvee_{t\in E}p_{t}$.
@@
Let $B$ be a $\mathrm{C}^{*}$-algebra. By considering the usual universal representation of $B$, we find a Hilbert space $H_{u}$ such that $B^{**}\subset B(H_{u})$. As in 2.4.1 in [pedersen1979book] let $(B_{\text{sa}})^{m}\subset B^{**}$ denote the set of elements in $B^{**}$ that can be obtained as the increasing limit in the SOT of $B(H_{u})$ of a self-adjoint net from $B$. A self-adjoint element $x\in B^{**}$ is said to be universally measurable if, for every $\varepsilon>0$ and state $\varphi$ on $B$, there are elements $y,z\in(B_{\text{sa}})^{m}$ such that $y\leq x\leq z$ and $\varphi(y+z)<\varepsilon$.
Projections in $B^{**}$ that are either open or closed are universally measurable [pedersen1979book, Proposition 3.11.9]. More generally, the class of universally measurable projections also contains the Borel projections in the sense of [combes1970]; see [brown2014] for more detail.
@@
Let $z_{\operatorname{at}}\in B^{**}$ denote the supremum of all minimal projections in $B^{**}$. It is well known that $z_{\operatorname{at}}$ is central: indeed, it coincides with the support projection for the unique weak-* continuous extension to $B^{**}$ of the atomic representation of $B$ [pedersen1979book, Paragraph 4.3.7]. In particular, by virtue of [pedersen1979book, Lemma 4.3.8], we see that there is a collection $\{H_{\lambda}:\lambda\in\Lambda\}$ of Hilbert spaces such that $z_{\operatorname{at}}B^{**}\cong\prod_{\lambda\in\Lambda}B(H_{\lambda})$. Consequently, if $p\in B^{**}$ is a projection, we have
@@
Let $p,e\in B^{**}$ be universally measurable projections with $e\geq pz_{\operatorname{at}}$. Since $z_{\operatorname{at}}$ is central, we find $(e-p)z_{\operatorname{at}}\geq 0$. In turn, because epe-p is universally measurable [pedersen1979book, Proposition 4.3.13], we infer that $e\geq p$ by virtue of [pedersen1979book, Theorem 4.3.15]. ∎
@@ 4. The non-commutative Varopoulos phenomenon
Let $B$ be a $\mathrm{C}^{*}$-algebra, let $q\in B^{**}$ be a closed projection and let $n\geq 1$. Let $R\in{\mathbb{M}}_{1,n}(qB^{**})$ and $C\in{\mathbb{M}}_{n,1}(B^{**}q)$. Fix $\Delta\in{\mathbb{M}}_{n}(B)$, a functional $\varphi\in B^{*}$ and a number $\varepsilon>0$. Then, there are $S\in{\mathbb{M}}_{1,n}(qB)$, $T\in{\mathbb{M}}_{n,1}(Bq)$ such that $\|S\|\leq\|R\|$$\|T\|\leq\|C\|$ and $|\varphi(R\Delta C-S\Delta T)|<\varepsilon.$
@@
By [hay2007, Proposition 3.1], it follows that $\Phi$ is a completely isometric isomorphism. Hence, $\Phi^{**}:(B/J)^{**}\to(qB)^{**}$ is a completely isometric weak-* homeomorphism. In turn, by [BLM2004, Paragraph 1.4.4], there is a completely isometric weak-* homeomorphism $\Psi:(B/J)^{**}\to B^{**}/J^{\perp\perp}$ such that
@@
Next, since $q$ is closed, and is the support projection of $J$ [akemann1969], we obtain that $J^{\perp\perp}=(1-q)B^{**}$. Hence, the map $\Theta:B^{**}/J^{\perp\perp}\to qB^{**}$ defined as
@@
is also a completely isometric weak-* homeomorphism. Therefore, we obtain a completely isometric weak-* homeomorphism $\Omega=\Theta\circ\Psi\circ(\Phi^{**})^{-1}:(qB)^{**}\to qB^{**}$ acting as the inclusion on $qB$.
@@
acting as the inclusion on ${\mathbb{M}}_{1,n}(qB)$. Then, the map $\Xi=\Omega^{(n)}\circ\Gamma:{\mathbb{M}}_{1,n}(qB)^{**}\to{\mathbb{M}}_{1,n}(qB^{**})$ is an isometric weak-* homeomorphism acting as the inclusion on ${\mathbb{M}}_{1,n}(qB)$. Consider the weak-* continuous linear functional $\psi$ on ${\mathbb{M}}_{1,n}(qB)^{**}$ defined as
@@
By Goldstine’s lemma (namely that the unit ball of a Banach space is weak* dense in the ball of the bidual), we can find $S_{0}\in{\mathbb{M}}_{1,n}(qB)$ with $\|S_{0}\|\leq\|\Xi^{-1}(R)\|=\|R\|$ such that
@@
To find $T$, we argue in a similar fashion. First, proceed as above to find an isometric weak-* homeomorphism $\Xi^{\prime}:{\mathbb{M}}_{n,1}(Bq)^{**}\to{\mathbb{M}}_{n,1}(B^{**}q)$ acting as the inclusion on ${\mathbb{M}}_{n,1}(Bq)$. Consider the weak-* continuous linear functional $\psi^{\prime}$ on ${\mathbb{M}}_{n,1}(Bq)^{**}$ defined as
@@
Let $B$ be a $\mathrm{C}^{*}$-algebra and let $A\subset B$ be a subalgebra. A projection $e\in B^{**}$ is said to be $A$-invariant if $ae=eae$ for each $a\in A$. Further, $e$ is said to be $A$-coinvariant if 1e1-e is $A$-invariant. We can now state and prove one of the main results of the paper.
@@
Let $0<\varepsilon<1$ and let $\varphi\in A^{\perp}$. We may assume without loss of generality that $\|\varphi\|=1$, so that $|\varphi|$ is a state. Throughout, we put $\tau=|\varphi|$. There is a partial isometry $y\in B^{**}$ such that $\tau=\varphi(\cdot y)$ and $\varphi=\tau(\cdot y^{*})$. By Goldstine’s lemma, there is $t\in B$ such that $\|t\|\leq 1$ and $|\varphi(qy-qt)|<\varepsilon.$ Because $q$ is a complete $(A,\kappa)$-interpolation projection, there is $x\in A$ such that $qx=qt$ and $\|x\|\leq\kappa$. In particular, we have
(5) $\tau(q)=|\varphi(qy)|<\varepsilon+|\varphi(qx)|.$
@@ 4.1. Atomic parts and double annihilators
Our class of examples will involve tensor products. We recall some relevant facts here, since several preliminaries are needed. Let $B$ and $D$ be $\mathrm{C}^{*}$-algebras such that $B$ is nuclear. It follows from [AB1980, Theorem 3.1] that there is an injective *-homomorphism $\sigma:B^{**}\otimes_{\min}D^{**}\to(B\otimes_{\min}D)^{**}$ with the following properties.
@@
    Let $\sigma_{B}:B^{**}\to(B\otimes_{\min}D)^{**}$ and $\sigma_{D}:D^{**}\to(B\otimes_{\min}D)^{**}$ be the restriction *-homomorphisms arising from $\sigma$ (see for instance [BO2008, Theorem 3.2.6]). Then, $\sigma_{B}$ and $\sigma_{D}$ are both weak-* continuous.
@@
We first assume that $r$ is open. Let $(d_{j})$ be a net of positive contractions in $D$ increasing to $r$ in the weak-* topology of $D^{**}$. Then $(\sigma_{D}(d_{j}))$ increases to $\sigma_{D}(r)$ in the weak-* topology of $(B\otimes_{\min}D)^{**}$. In the universal representation of $B\otimes_{\min}D$, this implies that in fact $(\sigma_{D}(d_{j}))$ increases to $\sigma_{D}(r)$ in the strong operator topology. On the other hand, by Goldstine’s lemma there is a net $(a_{i})$ of contractions in $A$ converging to $q$ in the weak-* topology of $B^{**}$. We see that the net $(\sigma_{B}(a_{i})\sigma_{D}(d_{j}))_{(i,j)}$ converges to $\sigma(q\otimes r)$ in the weak-* topology of $(B\otimes_{\min}D)^{**}$, being the product of a bounded weak-* converging net and a bounded strongly converging net. Each $\sigma_{B}(a_{i})\sigma_{D}(d_{j})=\sigma(a_{i}\otimes d_{j})$ lies in the image of $A\otimes D$ inside of $(B\otimes_{\min}D)^{**}$, and therefore $q\otimes r\in(A\otimes_{\min}D)^{\perp\perp}$.
@@
Let $q\in(\mathrm{C}(X)\otimes_{\min}D)^{**}$ be a minimal projection. By Lemma 4.5, there are minimal projections $e\in\mathrm{C}(X)^{**}$ and $f\in D^{**}$ such that $q=e\otimes f$. As is well-known, $e$ is simply the characteristic function of a point $x\in X$. By [gamelin1969, Theorem II.11.3], we see that $x\in X$ is an $A$-peak point, since it lies in the Choquet boundary. Hence $e\in A^{\perp\perp}$, so by Lemma 4.4, we conclude that $q\in(A\otimes_{\min}D)^{\perp\perp}$. This proves the first statement. The second statement immediately follows from the first along with (2). ∎
@@ 4.2. Recovering Varopoulos’ null projections
    there is a net $(b_{i})$ of positive contractions in $B$ commuting with $A$ and decreasing to $q$ in the weak-* topology of $B^{**}$,
@@ 5.1. Adapted projections
Let $B$ be a $\mathrm{C}^{*}$-algebra and let $A\subset B$ be a subalgebra. We say that a projection $e\in B^{**}$ is $A$-adapted if there is a net $(e_{i})$ of closed $A$-coinvariant projections in $B^{**}$ increasing to $e$ in the weak-* topology of $B^{**}$. The linchpin technical result of this section is the following refinement of Theorem 4.3.
@@
Let $B$ be a unital $\mathrm{C}^{*}$-algebra and let $A\subset B$ be a unital norm-closed subalgebra. Assume that there is a central closed projection $z\in B^{**}$ such that $B^{**}z$ is commutative and 1z1-z is $A$-adapted. Let $q\in B^{**}$ be a projection with the following properties:
@@
Further, because 1z1-z is $A$-adapted, there is an $A$-coinvariant closed projection $r\leq 1-z$ such that
@@
Because $z$ is closed, so is z(1u)z(1-u) [akemann1969, Proposition II.5]. Moreover, we have $r\leq 1-z$ so that $r^{\prime}=r+z(1-u)$ is closed by [akemann1969, Theorem II.7]. The projection
@@
Observe next that since $B^{**}z$ is commutative, $q=qz$ is $A$-coinvariant and $uz$ is $A$-invariant. On the other hand, 1z1-z is central and $r$ is $A$-coinvariant, so that (1z)r(1-z)-r is $A$-invariant. It follows that $u^{\prime}$ is $A$-invariant as well. By virtue of Theorem 4.3, we see that
@@ 5.2. Hilbert function spaces and their multipliers
A multiplier is a function $\varphi:{\mathbb{B}}_{d}\to{\mathbb{C}}$ satisfying $\varphi f\in{\mathcal{H}}$ for every $f\in{\mathcal{H}}$. The corresponding multiplication operator $M_{\varphi}:{\mathcal{H}}\to{\mathcal{H}}$ is then bounded. The space of all multiplication operators induced by a multiplier is denoted by ${\mathcal{M}}({\mathcal{H}})$, and it is weak-* closed subalgebra of $B({\mathcal{H}})$. If $\varphi$ is a multiplier, then
@@
For each finite subset $F\subset{\mathbb{B}}_{d}$, let $e_{F}\in B({\mathcal{H}})$ denote the projection onto the subspace spanned by $\{k_{\lambda}:\lambda\in F\}$. Since the set $\{k_{\lambda}:\lambda\in{\mathbb{B}}_{d}\}$ spans a norm-dense subset of $\mathcal{H}$}, the net $(e_{F})$ increases to the identity on $\mathcal{H}$} in the weak-* topology of $B({\mathcal{H}})$. Further, each $e_{F}$ is compact on $\mathcal{H}$}, and it is ${\mathcal{A}}({\mathcal{H}})$-coinvariant by virtue of (16). Next, by [CT2023Henkin, Lemma 5.5], there is a weak-* homeomorphic *-isomorphism $\Omega:{\mathcal{T}}({\mathcal{H}})^{**}(1-z_{K})\to B({\mathcal{H}})$ such that $\Omega(t(1-z_{K}))=t$ for every $t\in{\mathcal{T}}({\mathcal{H}})$. Since $e_{F}$ is compact, it lies in ${\mathcal{T}}({\mathcal{H}})$ and we see that $\Omega^{-1}(e_{F})=e_{F}(1-z_{K})=e_{F}$. It follows that $(\Omega^{-1}(e_{F}))$ is an increasing net of ${\mathcal{A}}({\mathcal{H}})$-coinvariant closed projections in ${\mathcal{T}}({\mathcal{H}})^{**}$ increasing to $1-z_{K}$ in the weak-* topology of ${\mathcal{T}}({\mathcal{H}})^{**}$. ∎
Let $\gamma:{\mathcal{T}}({\mathcal{H}})\to{\mathcal{T}}({\mathcal{H}})/K$ denote the quotient map. There is a natural *-isomorphism $\tau:{\mathcal{T}}({\mathcal{H}})/K\to\mathrm{C}({\mathbb{S}}_{d})$ such that
@@
See e.g. [GHX04]. In particular, ${\mathcal{T}}({\mathcal{H}})$ is nuclear [BO2008, Exercise 3.8.1]. The restriction of $(\tau\circ\gamma)^{**}$ gives a weak-* homeomorphic *-isomorphism $\Theta:{\mathcal{T}}({\mathcal{H}})^{**}z_{K}\to\mathrm{C}({\mathbb{S}}_{d})^{**}$. We now give a description of the closed projections in ${\mathcal{T}}({\mathcal{H}})^{**}$ dominated by $z_{K}$.
@@
(i) Let $(t_{i})$ be a net of positive contractions in ${\mathcal{T}}({\mathcal{H}})$ decreasing to $q$ in the weak-* topology of ${\mathcal{T}}({\mathcal{H}})^{**}$. Since $\Theta$ is a weak-* continuous *-homomorphism with $\Theta(z_{K})=1$, we have that $\Theta(t_{i}z_{K})=(\tau\circ\gamma)^{**}(t_{i})=(\tau\circ\gamma)(t_{i})\in\mathrm{C}({\mathbb{S}}_{d})$. It follows that $(\Theta(t_{i}z_{K}))$ is a net of positive contractions in $\mathrm{C}({\mathbb{S}}_{d})$ decreasing to $\Theta(q)$ in the weak-* topology of $\mathrm{C}({\mathbb{S}}_{d})^{**}$. In other words, $\Theta(q)$ is closed in $\mathrm{C}({\mathbb{S}}_{d})^{**}$, so there is a closed subset $E\subset{\mathbb{S}}_{d}$ such that $\Theta(q)=\chi_{E}$.
@@
$\chi_{E}\,\{\varphi:M_{\varphi}\in{\mathcal{A}}({\mathcal{H}})\}=\chi_{E}\,\mathrm{C}({\mathbb{S}}_{d})$
@@
    $q{\mathcal{A}}({\mathcal{H}})=q{\mathcal{T}}({\mathcal{H}})$.
@@
By Lemma 5.3, there is a closed subset $E\subset{\mathbb{S}}_{d}$ such that $\Theta(q)=\chi_{E}$. Since $\Theta$ is a *-isomorphism such that